hallm:

1. What is the product of and ? Write your answer in standard form. (a) Show your work. (b) Is the product of and equal to the product of and ? Explain your answer.

Shadow:

Identify what you need to find the product of. It's blank.

hallm:

its what is the product of 1/2x - 1/4 and 5x^2 - 2x + 6

Shadow:

\[(\frac{ 1 }{ 2 }x - \frac{ 1 }{ 4 }) \times (5x^2 - 2x + 6)\] This?

hallm:

and b is the product of 1/2x - 1/4 and 5x^2 - 2x equal to the product of 1/4x - 1/2 and 5x^2 - 2x + 6 and yes thats it

Shadow:

For a, \[\frac{ 1 }{ 2 }x \times (5x^2 -2x + 6)\] Multiply 1/2x to each term \[-\frac{ 1 }{ 4 } \times (5x^2 -2x + 6)\] Multiply -1/4 to each term Then add up both of the results. The sum of each.

Shadow:

For b you can check your work by doing the exact same method.

Shadow:

@hallm Does that make sense?

hallm:

yes thank you